How do i write a quadratic function in standard formwith the zeros; -4 and -4
Found 2 solutions by mananth, stanbon:
Answer by mananth(16949)

(x+4)(x+4)
x^2+8x+16

Answer by stanbon(75887)

How do i write a quadratic function in standard formwith the zeros; -4 and -4
----
If -4 is a zero, x+4 is a factor.
---
f(x) = (x+4)(x+4)
---
f(x) = (x+4)^2
---
f(x) = x^2+8x+16
====================
cheers,
Stan H.